# MV-VTON: Multi-View Virtual Try-On with Diffusion Models

### v2

@@ Abstract.
The goal of image-based virtual try-on is to generate an image of the target person naturally wearing the given clothing. However, most existing methods solely focus on the frontal try-on using the frontal clothing. When the views of the clothing and person are significantly inconsistent, particularly when the person’s view is non-frontal, the results are unsatisfactory. To address this challenge, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to reconstruct the dressing results of a person from multiple views using the given clothes. On the one hand, given that single-view clothes provide insufficient information for MV-VTON, we instead employ two images, i.e., the frontal and back views of the clothing, to encompass the complete view as much as possible. On the other hand, the diffusion models that have demonstrated superior abilities are adopted to perform our MV-VTON. In particular, we propose a view-adaptive selection method where hard-selection and soft-selection are applied to the global and local clothing feature extraction, respectively. This ensures that the clothing features are roughly fit to the person’s view. Subsequently, we suggest a joint attention block to align and fuse clothing features with person features. Additionally, we collect a MV-VTON dataset, i.e., Multi-View Garment (MVG), in which each person has multiple photos with diverse views and poses. Experiments show that the proposed method not only achieves state-of-the-art results on MV-VTON task using our MVG dataset, but also has superiority on frontal-view virtual try-on task using VITON-HD and DressCode datasets. Codes and datasets will be publicly released at https://github.com/hywang2002/MV-VTON.
Virtual try-on, Multi-view, Diffusion models
ccs: Computing methodologies Computer vision
1. Introduction
@@
Virtual Try-On (VTON) is a classic yet intriguing technology. It can be applied in the field of fashion, and clothes online shopping to improve user experience. VTON aims to render the visual effect of a person wearing a specified garment. The emphasis of this technology lies in reconstructing a realistic image that faithfully preserves personal attributes (e.g., appearance and pose) and accurately represents clothing shape and details.
@@
Most early VTON methods (Choi et al., 2021; Lee et al., 2022; Xie et al., 2023; Bai et al., 2022; Han et al., 2019; He et al., 2022; Li et al., 2021; Yang et al., 2020) are based on generative adversarial networks (Goodfellow et al., 2020) (GANs). They generally align the clothing to the person’s pose through a warping network, and then employ a generator to fuse the warped clothing with the person. However, it poses a challenge to ensure that the warped clothing fits the target person’s pose, and inaccurate clothes features will easily lead to distortion results. Recently, the diffusion models (Rombach et al., 2022; Zhang et al., 2023; Huang et al., 2023c; Nichol and Dhariwal, 2021) have made remarkable strides in the field of image generation (Li et al., 2023; Ruiz et al., 2023; Huang et al., 2023a; Xu et al., 2023; Huang et al., 2023b). Leveraging its potent generative capabilities, some researchers (Morelli et al., 2023; Gou et al., 2023; Kim et al., 2023; Zhu et al., 2023; Xu et al., 2024) have integrated it into virtual try-on endeavors, building upon previous work and achieving commendable results. For example, LaDI-VTON (Morelli et al., 2023) encodes the garment through textual inversion as the conditional input for Stable Diffusion (Rombach et al., 2022) backbone, and introduces a residual connection between the VAE (Kingma and Welling, 2013) encoder and decoder to preserve person’s face and hands. DCI-VTON (Gou et al., 2023) attaches the pre-warped garment to the inpainting area, providing a priori information for the diffusion model. StableVITON (Kim et al., 2023) introduces an additional encoder and attention blocks to deform the clothing.
@@
Although VTON has made great progress, most existing methods focus on taking the frontal garment to perform the frontal try-on. In practical applications, such as online shopping for clothes, customers may expect to obtain the dressing effect on multiple views (e.g., side or back). In this case, the pose of the garment may be seriously inconsistent with the person’s posture, and the single-view clothing may not be enough to provide complete try-on information. Thus, these methods will easily generate results with poorly deformed clothing, and lead to the loss of high-frequency details such as texts, patterns, and other textures on clothing.
@@
To address the issue, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to faithfully reconstruct the appearance and attire of a person from various views. MV-VTON should not only address the challenge of natural warping on the clothing’s front but also tackle the issue of inconsistency between the poses of the given clothing and target person. For example, for hooded sweaters, which may exhibit significant differences between frontal and back styles, MV-VTON should display them from various views to provide a more comprehensive fitting result. Thus, only providing single clothing can’t meet the needs of dressing up. Instead, we utilize both the frontal and back views of the clothing, which covers approximately the complete view with as few images as possible.
@@
Given the frontal and back clothing, we utilize the popular diffusion method to achieve MV-VTON. It is natural but doesn’t work well to simply concatenate two pieces of clothing together as conditions of diffusion models, as it is difficult for the model to learn how to assign two-view clothes to a person, especially when the person is sideways. Instead, we propose a view-adaptive selection mechanism, which picks appropriate features of two-view clothes based on the posture information of the person and clothes. Therein, the hard-selection module chooses one of the two clothes for global feature extraction, and the soft-selection module modulates the local features of two clothes. We utilize CLIP (Radford et al., 2021) and an additional multi-scale encoder to extract the global and local clothing features, respectively. Moreover, to enhance the preservation of high-frequency details in clothing, we present the joint attention block. It independently aligns global and local features with the person features, and selectively fuses them to refine the local clothing details while preserving global semantic information.
@@
Furthermore, we collect a multi-view virtual try-on dataset, named Multi-View Garment (MVG). It contains thousands of samples, and each sample contains 5 images under different views and poses. We conduct extensive experiments not only on MV-VTON task using the MVG dataset, but also on the frontal-view VTON task using VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. The results demonstrate that our method outperforms existing methods on both tasks, quantitatively and qualitatively.
@@
  • We propose the joint attention block to align the global and local features of selected clothing with the person ones, and fuse them.
@@
2. Related Work
2.1. GAN-Based Virtual Try-On
@@
Most of the previous methods are aimed at the frontal-view VTON task. To reconstruct realistic results, existing methods based on generative adversarial networks (GAN) (Goodfellow et al., 2020) are typically divided into two steps. Firstly, the frontal-view clothing is deformed to initially align with the target person’s pose. Afterward, the warped clothing and target person are fused through a GAN-based generator. In the warping step, some methods (Morelli et al., 2023; Yang et al., 2020; Ge et al., 2021a; Wang et al., 2018) use TPS transformation to deform the frontal-view flattened clothing, and others (Lee et al., 2022; Ge et al., 2021b; Xie et al., 2023) predict the global and local optical flow required for clothing deformation. However, when the clothing possesses intricate high-frequency details and the person’s pose is complex, the effectiveness of clothing deformation is often significantly diminished. Moreover, GAN-based generators generally encounter challenges in convergence and are highly susceptible to mode collapse (Miyato et al., 2018), leading to noticeable artifacts at the junction between warped clothing and the target person in the final results. In addition, previous multi-pose virtual try-on methods (Dong et al., 2019; Wang et al., 2020; Yu et al., 2023) can change the person’s pose, but are also limited by GAN-based generator and insufficient clothing information. In this paper, we leverage the potent generative capability of a large-scale pre-trained diffusion model to implicitly deform the clothing from frontal and back views. Simultaneously, we employ the view-adaptive selection to choose the clothing features from the two views.
2.2. Diffusion-Based Virtual Try-On
@@
Thanks to the rapid advancement of diffusion models, recent works have sought to utilize the generative prior of large-scale pre-trained diffusion models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022; Yang et al., 2023) to tackle frontal-view virtual try-on tasks. TryOnDiffusion (Zhu et al., 2023) introduces two U-Nets to encode target person and frontal-view clothing images respectively, and interacts with the features of the two branches through the cross-attention mechanism. LaDI-VTON (Morelli et al., 2023) encodes the frontal-view clothing image through textual inversion (Gal et al., 2022; Wei et al., 2023) and serves as the conditional input of backbone. DCI-VTON (Gou et al., 2023) first conducts an initial deformation of frontal-view clothing by incorporating a pre-trained wrapping network (Ge et al., 2021b). Subsequently, it attaches the deformed clothing to the target person image and feeds it into the diffusion model. While their frontal-view virtual try-on results seem more natural compared to GAN-based methods, they face difficulties in preserving high-frequency details due to the loss of details from the CLIP image encoder (Radford et al., 2021). To address this problem, StableVITON (Kim et al., 2023) attempts to introduce an additional encoder (Zhang et al., 2023) to encode the local features of frontal-view clothing, and align the obtained clothing features through the zero cross-attention block. However, StableVITON (Kim et al., 2023) aligns the global features and local features of clothing separately without considering direct interaction between the two, making it difficult to refine the details while preserving global semantic information. Therefore, we introduce the joint attention blocks, building upon the foundation of maintaining two encoders to extract the global and local features of clothing, respectively. We first align the global and local features to the target person, then selectively fuse the two to enhance the preservation of local texture and global semantic information in the original clothing.
3. method
While most existing virtual try-on methods are designed solely for frontal-view scenarios, we present a novel approach to handle both frontal-view and multi-view virtual try-on tasks, along with a multi-view virtual try-on dataset MVG comprising try-on images captured from five different views. Examples of our dataset are shown in Figure 2(b). Formally, given a person image $x\in R^{H\times W\times 3}$ in a frontal, side, or back view, along with a frontal view clothing $c_{f}\in R^{H\times W\times 3}$ and a back view clothing $c_{b}\in R^{H\times W\times 3}$, our goal is to generate the result of the person wearing the clothing in its view. Considering the substantial differences between the front and back of most clothing items, compared to previous frontal-view virtual try-on methods, another challenge we face is how to make informed decisions regarding the two provided clothing images based on the target person’s pose, ensuring a natural try-on result across multiple views.
In this work, we use an image inpainting diffusion model (Yang et al., 2023) as the baseline of our method, it concatenates the noisy image $z_{t}$, the encoded clothing-agnostic image $\mathcal{E}(a)$, and the resized clothing-agnostic mask $m$ in the channel dimension and feeds them into the backbone as spatial input. Therein, $a$ is the masked person image $x$ using the inpainting mask $M\in R^{H\times W\times 1}$. In addition, we use an existing method to initially deform the clothing and paste it on $a$. $m$ is the result of resizing $M$ to latent space and $\mathcal{E}$ represents VAE encoder (Kingma and Welling, 2013) in latent diffusion model (Rombach et al., 2022). While utilizing CLIP image encoder (Radford et al., 2021) to encode clothing as the global condition of the diffusion model, we also introduce an additional encoder (Yang et al., 2023; Zhang et al., 2023) to encode clothing to provide more refined local conditions. Since both the frontal and back view clothing need to be encoded, directly sending both into the backbone as conditions may result in confusion of clothing features. To alleviate this problem, we propose a view-adaptive selection mechanism. Based on the similarity between poses of the person and two clothes, it conducts hard-selection when extracting global features and soft-selection when extracting local features. To enhance high-frequency details in global features using local ones, while preserving semantic information in clothing, we introduce a joint attention block. It first independently aligns global and local features to the person ones and then selectively fuses them. Figure 3(a) depicts an overview of our proposed method.
Figure 2. Comparison between previous datasets and our proposed MVG dataset. (a) is the datasets used by the previous work, which only have clothing and person in the frontal-view. In contrast, each sample in our dataset (b) offers images from five different views. By utilizing the first and fifth images to provide frontal and back view clothing, we can reconstruct multi-view virtual try-on results.
[…]
[Figure 2. Fully described in the text.]
Figure 3. (a) Overview of MV-VTON. It encodes frontal and back view clothing into global features using the CLIP image encoder (Radford et al., 2021) and extracts multi-scale local features through backbone encoder’s trainable copy $\mathcal{E}_{l}$. Both features act as conditional inputs for the decoder of backbone. Besides, both features are selectively extracted through view-adaptive selection mechanism. (b) Soft-selection modulates the clothing features on frontal and back view, respectively, based on the similarity between the clothing’s pose and the person’s pose. Then the features from both views are concatenated in the channel dimension.
[…]
[Figure 3. Fully described in the text.]
3.1. Preliminaries for Diffusion Models
@@
Diffusion Models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022) have demonstrated strong capabilities in visual generation, which transforms a Gaussian distribution into a target distribution by iterative denoising. In particular, Stable Diffusion (Rombach et al., 2022) is a widely used generative diffusion model, which consists of a CLIP text encoder  (Radford et al., 2021) $\mathcal{E}_{T}$, a VAE (Kingma and Welling, 2013) encoder $\mathcal{E}$ as well as decoder $\mathcal{D}$, and a time-conditional denoising model $\epsilon_{\theta}$. The text encoder $\mathcal{E}_{T}$ encodes the input text prompt $y$ as conditional input. The VAE encoder $\mathcal{E}$ compresses the input image $I$ into latent space to get the latent variable $z_{0}=\mathcal{E}(I)$. In contrast, the VAE decoder $\mathcal{D}$ decodes the output of backbone from latent space to pixel space. Through the VAE encoder $\mathcal{E}$, at an arbitrary time step t, the forward process is performed:
(1) $\alpha:={\textstyle\prod_{s=1}^{t}(1-\beta_{s})}~{},\indent z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$
@@
where $\epsilon\sim\mathcal{N}(0,1)$ is the random Gaussian noise and $\beta$ is a predefined variance schedule. The training objective of the diffusion model is to acquire a noise prediction network that minimizes the disparity between the predicted noise and the noise added to ground truth. The loss function can be defined as,
(2) $\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),y,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{T}(y))\right\|_{2}^{2}],$
@@
where $t$ is the time step, and $z_{t}$ represents the encoded image $\mathcal{E}(I)$ with random Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$ added.
@@
In our work, we use an exemplar-based inpainting model (Yang et al., 2023) as a backbone, which employs an image $c$ rather than texts as the prompt and then encode $c$ by the image encoder $\mathcal{E}_{I}$ of CLIP (Radford et al., 2021). Thus, the loss function in Eq. (2) can be modified as,
(3) $\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),c,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{I}(c))\right\|_{2}^{2}].$
3.2. View-Adaptive Selection
@@
For multi-view virtual try-on task, given the substantial differences between the frontal and back view, as illustrated in Figure 2(b), it’s imperative to extract and assign the features of frontal and back view clothing for the person tendentiously. Actually, based on the pose of the target person, we can determine which view’s clothing should be given more attention during the try-on process. For example, if the target pose resembles the pose in the fourth column of Figure 2(b), it’s evident that we should rely more on the characteristics of the back view clothing to generate the try-on result. Specifically, we propose a view-adaptive selection mechanism to achieve this purpose, including hard- and soft-selection.
@@
Hard-Selection for Global Clothing Features. The image encoder of CLIP (Radford et al., 2021) is used to extract global features of clothing. Before this process, we perform hard-selection on the clothing images under the frontal and back view based on the similarity between the garments’ pose and the person’s pose. It means that we only select one piece of clothing that is closest to the person’s pose as the input of the image encoder, since it is enough to cover global semantic information. When generating pre-warped clothing in $\mathcal{E}(a)$, the selection is also performed. Implementation details of hard-selection can be found in the supplementary material.
@@
Soft-Selection for Local Clothing Features. The multi-scale local features of frontal and back view clothing are extracted through an additional encoder $\mathcal{E}_{l}$. Considering that relying solely on the clothing from the frontal or back view may not be sufficient when reconstructing the try-on results under certain specific scenes, such as the third column shown in Figure 2(b). In these cases, it may be necessary to incorporate clothing features from both the frontal and back views. However, simply combining the two may lead to confusion of features. Instead, we introduce soft-selection block to modulate their local features, respectively, as shown in Figure 3(b). Firstly, the person’s pose $p_{h}$, frontal-view clothing’s pose $p_{f}$ and back view clothing’s pose $p_{b}$ are encoded by the pose encoder $\mathcal{E}_{p}$ to obtain their respective features $\mathcal{E}_{p}(p_{h})$, $\mathcal{E}_{p}(p_{f})$ and $\mathcal{E}_{p}(p_{b})$. Details of the pose encoder can be found in the supplementary material. When processing frontal-view clothing, in $i$-th soft-selection block, we map $\mathcal{E}_{p}(p_{h})$ and $\mathcal{E}_{p}(p_{f})$ to $P^{i}_{h}$ and $P^{i}_{f}$ through a linear layer with weights $W_{h}^{i}$ and $W_{f}^{i}$, respectively. For the features of frontal-view clothing $c_{f}^{i}$, it will be mapped to $C^{i}_{f}$ through a linear layer with weights $W_{c}^{i}$. Then we calculate the similarity between the person’s pose and frontal-view clothing’s pose to get the selection weights of frontal-view clothing, i.e.,
(4) $weights=softmax(\frac{P_{h}^{i}(P_{f}^{i})^{T}}{\sqrt{d}}),$
@@
where $weights$ represents the selection weights of frontal-view clothing, and $d$ represents the dimension of these matrices. Assuming that the person’s pose is biased towards the front, as depicted in the second column of Figure 2(b), the similarity between the person’s pose and the front view clothing’s pose will be higher. Consequently, the corresponding clothing features will be enhanced by $weights$, and vice versa.
The features of back view clothing $c_{b}^{i}$ undergo similar processing. Consequently, one of the features $c_{f}^{i}$ and $c_{b}^{i}$ in the two views will be enhanced while the other will be weakened. Finally, the two selected clothing features are concatenated along the channel dimension as the local condition $c_{l}^{i}$ of backbone.
Figure 4. Overview of the proposed joint attention block.
[…]
[Figure 4. Fully described in the text.]
3.3. Joint Attention Block
@@
Global clothing features $c_{g}$ provide identical conditions for blocks at each scale of U-Net, and multi-scale local clothing features $c_{l}$ allow for reconstructing more accurate details. We present a joint attention block to align $c_{g}$ and $c_{l}$ with the current person features, as shown in Figure 4. Additionally, to retain most of the semantic information in global features $c_{g}$, we use local features $c_{l}$ to refine some lost and erroneous detailed texture information in $c_{g}$ by selective fusion.

### v4

@@ Abstract
The goal of image-based virtual try-on is to generate an image of the target person naturally wearing the given clothing. However, existing methods solely focus on the frontal try-on using the frontal clothing. When the views of the clothing and person are significantly inconsistent, particularly when the person’s view is non-frontal, the results are unsatisfactory. To address this challenge, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to reconstruct the dressing results from multiple views using the given clothes. Given that single-view clothes provide insufficient information for MV-VTON, we instead employ two images, i.e., the frontal and back views of the clothing, to encompass the complete view as much as possible. Moreover, we adopt diffusion models that have demonstrated superior abilities to perform our MV-VTON. In particular, we propose a view-adaptive selection method where hard-selection and soft-selection are applied to the global and local clothing feature extraction, respectively. This ensures that the clothing features are roughly fit to the person’s view. Subsequently, we suggest joint attention blocks to align and fuse clothing features with person features. Additionally, we collect a MV-VTON dataset MVG, in which each person has multiple photos with diverse views and poses. Experiments show that the proposed method not only achieves state-of-the-art results on MV-VTON task using our MVG dataset, but also has superiority on frontal-view virtual try-on task using VITON-HD and DressCode datasets.
Codehttps://github.com/hywang2002/MV-VTON
Figure 1: Motivation of this work. Previous VTON methods, e.g., StableVITON (Kim et al. 2023) can only be used for the frontal-view person, and fail when facing the person with multiple views. Our MV-VTON can faithfully present the try-on results for a person with various views.
Introduction
@@
Virtual Try-On (VTON) is a classic yet intriguing technology. It can be applied in the field of fashion and clothes online shopping to improve user experience. VTON aims to render the visual effect of a person wearing a specified garment. The emphasis of this technology lies in reconstructing a realistic image that faithfully preserves personal attributes and accurately represents clothing shape and details.
@@
Early VTON methods (Lee et al. 2022; Xie et al. 2023; Bai et al. 2022; He, Song, and Xiang 2022) are based on generative adversarial networks (Goodfellow et al. 2020) (GANs). They generally align the clothing to the person’s pose, and then employ a generator to fuse the warped clothing with the person. However, it poses a challenge to ensure that the warped clothing fits the target person’s pose, and inaccurate clothes features will easily lead to distortion results. Recently, diffusion models (Rombach et al. 2022) have made remarkable strides in the field of image generation (Ruiz et al. 2023). Leveraging its potent generative capabilities, some researchers (Morelli et al. 2023; Kim et al. 2023) have integrated it into virtual try-on fields, building upon previous work and achieving commendable results.
@@
Although VTON has made great progress, most existing methods focus on performing the frontal try-on. In practical applications, such as online shopping for clothes, customers may expect to obtain the dressing effect on multiple views (e.g., side or back). In this case, the pose of the garment may be seriously inconsistent with the person’s posture, and the single-view clothing may not be enough to provide complete try-on information. Thus, these methods will easily generate results with poorly deformed clothing, and lead to the loss of high-frequency details such as texts, patterns, and other textures on clothing, as shown in Figure 1.
@@
To address these issues, we introduce Multi-View Virtual Try-ON (MV-VTON), which aims to reconstruct the appearance and attire of a person from multiple views. For example, for clothing in Figure 1, which may exhibit significant differences between frontal and back styles, MV-VTON should be able to display try-on results in various views, including front, back, and side ones. Thus, providing single clothing can’t meet the needs of dressing up, as the clothing only has partial information. Instead, we utilize both the frontal and back views of the clothing, which covers approximately complete view with as few images as possible.
@@
Given the frontal and back clothing, we utilize the popular diffusion method to achieve MV-VTON. It is natural but doesn’t work well to simply concatenate two pieces of clothing together as conditions of diffusion models, as it is difficult for the model to learn how to assign two-view clothes to a person, especially when the person is sideways. Instead, we propose a view-adaptive selection mechanism, which picks appropriate features of two-view clothes based on the posture information of the person and clothes. Therein, the hard-selection module chooses one of the two clothes for global feature extraction, and the soft-selection module modulates the local features of two clothes. We utilize CLIP (Radford et al. 2021) and a multi-scale encoder to extract the global and local clothing features, respectively. Moreover, to enhance the preservation of high-frequency details in clothing, we present joint attention blocks. They independently align global and local features with the person features, and selectively fuse them to refine the local clothing details while preserving global semantic information.
@@
Furthermore, we collect a multi-view virtual try-on dataset, named Multi-View Garment (MVG). It contains thousands of samples, and each sample contains 5 images under different views and poses. We conduct extensive experiments not only on MV-VTON task using the MVG dataset, but also on the frontal-view VTON task using VITON-HD (Lee et al. 2022) and DressCode (Morelli et al. 2022) datasets. The results demonstrate that our method outperforms existing methods on both tasks.
@@
  • We propose joint attention blocks to align the global and local features of selected clothing with the person ones, and fuse them.
@@
Related Work
GAN-Based Virtual Try-On
@@
Existing methods are aimed at the frontal-view VTON task. To reconstruct realistic results, these methods based on generative adversarial networks (GAN) (Goodfellow et al. 2020) are typically divided into two steps. Firstly, the frontal-view clothing is deformed to align with the target person’s pose. Afterward, the warped clothing and target person are fused through a GAN-based generator. In the warping step, some methods (Yang et al. 2020; Ge et al. 2021a; Wang et al. 2018) use TPS transformation to deform the frontal-view clothing, and others (Lee et al. 2022; Ge et al. 2021b; Xie et al. 2023) predict the global and local optical flow required for clothing deformation. However, when the clothing possesses intricate high-frequency details and the person’s pose is complex, the effectiveness of clothing deformation is often diminished. Moreover, GAN-based generators generally encounter challenges in convergence and are highly susceptible to mode collapse (Miyato et al. 2018), leading to noticeable artifacts at the junction between warped clothing and the target person in the final results. In addition, previous multi-pose virtual try-on methods (Dong et al. 2019; Wang et al. 2020; Yu et al. 2023) can change the person’s pose, but are also limited by GAN-based generator and insufficient clothing information.
Diffusion-Based Virtual Try-On
Thanks to the rapid advancement of diffusion models, recent works have sought to utilize the generative prior of large-scale pre-trained diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Rombach et al. 2022; Yang et al. 2023) to tackle frontal-view virtual try-on tasks. TryOnDiffusion (Zhu et al. 2023) introduces two U-Nets to encode target person and frontal-view clothing images respectively, and interacts with the features of the two branches through the cross-attention mechanism.
@@
LaDI-VTON (Morelli et al. 2023) encodes the frontal-view clothing image through textual inversion (Gal et al. 2022; Wei et al. 2023) and serves as the conditional input of backbone. DCI-VTON (Gou et al. 2023) first conducts an initial deformation of frontal-view clothing by incorporating a pre-trained wrapping network (Ge et al. 2021b). Subsequently, it attaches the deformed clothing to the target person image and feeds it into the diffusion model. While their frontal-view virtual try-on results seem more natural compared to GAN-based methods, they face difficulties in preserving high-frequency details due to the loss of details from the CLIP image encoder (Radford et al. 2021). To address this problem, StableVITON (Kim et al. 2023) attempts to introduce an additional encoder (Zhang, Rao, and Agrawala 2023) to encode the features of frontal-view clothing, and align the obtained clothing features through the zero cross-attention block. However, due to the absence of adequate clothing priors, the generated results often struggle to remain faithful to the original clothing. Therefore, we introduce joint attention blocks to extract the global and local features of clothing, and employ the view-adaptive selection to choose the clothing features from the two views.
Method
Preliminaries for Diffusion Models
@@
Diffusion Models (Ho, Jain, and Abbeel 2020; Rombach et al. 2022) have demonstrated strong capabilities in visual generation, which transforms a Gaussian distribution into a target distribution by iterative denoising. In particular, Stable Diffusion (Rombach et al. 2022) is a widely used generative diffusion model, which consists of a CLIP text encoder $\mathcal{E}_{T}$, a VAE encoder $\mathcal{E}$ as well as decoder $\mathcal{D}$, and a time-conditional denoising model $\epsilon_{\theta}$. The text encoder $\mathcal{E}_{T}$ encodes the input text prompt $y$ as conditional input. The VAE encoder $\mathcal{E}$ compresses the input image $I$ into latent space to get the latent variable $z_{0}=\mathcal{E}(I)$. In contrast, the VAE decoder $\mathcal{D}$ decodes the output of backbone from latent space to pixel space. Through the VAE encoder $\mathcal{E}$, at an arbitrary time step t, the forward process is performed:
$\alpha:={\textstyle\prod_{s=1}^{t}(1-\beta_{s})}~{},\indent z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$ (1)
@@
where $\epsilon\sim\mathcal{N}(0,1)$ is the random Gaussian noise and $\beta$ is a predefined variance schedule. The training objective is to acquire a noise prediction network that minimizes the disparity between the predicted noise and the noise added to ground truth. The loss function can be defined as,
$\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),y,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{T}(y))\right\|_{2}^{2}],$ (2)
@@
where $z_{t}$ represents the encoded image $\mathcal{E}(I)$ with random Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$ added.
@@
In our work, we use an exemplar-based inpainting model (Yang et al. 2023) as a backbone, which employs an image $c$ rather than texts as the prompt and then encode $c$ by the image encoder $\mathcal{E}_{I}$ of CLIP. Thus, the loss function in Eq. (2) can be modified as,
$\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(I),c,\epsilon\sim\mathcal{N}(0,1),t}[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{E}_{I}(c))\right\|_{2}^{2}].$ (3)
Method Overview
While existing virtual try-on methods are designed solely for frontal-view scenarios, we present a novel approach to handle both frontal-view and multi-view virtual try-on tasks, along with a multi-view virtual try-on dataset MVG comprising try-on images captured from five different views. Examples of it are shown in Figure 2(b). Formally, given a person image $x$ in an arbitrary view, along with a frontal view clothing $c_{f}$ and a back view clothing $c_{b}$, our goal is to generate the result of the person wearing the clothing in its view. Considering the substantial differences between the front and back of most clothing, another challenge is to make informed decisions regarding the two provided clothing images based on the target person’s pose, ensuring a natural try-on result across multiple views.
In this work, we use an image inpainting diffusion model (Yang et al. 2023) as our backbone. Denote by $M$ the inpainting mask, and denote by $a$ the masked person image $x$. The model concatenates $z_{t}$ ($z_{0}=\mathcal{E}(x)$), the encoded clothing-agnostic image $\mathcal{E}(a)$, and the resized clothing-agnostic mask $m$ in the channel dimension, and feeds them into the backbone as spatial input. Besides, we use an existing method to pre-warp the clothing and paste it on $a$. While utilizing CLIP image encoder to encode clothing as the global condition of the diffusion model, we also introduce an additional encoder (Zhang, Rao, and Agrawala 2023) to encode clothing to provide more refined local conditions. Since both the frontal and back view clothing need to be encoded, directly sending both into the backbone as conditions may result in confusion of clothing features. To alleviate this problem, we propose a view-adaptive selection mechanism. Based on the similarity between the poses of the person and two clothes, it conducts hard-selection when extracting global features and soft-selection when extracting local features. To preserve semantic information in clothing and enhance high-frequency details in global features using local ones, we introduce joint attention blocks. They first independently align global and local features to the person ones and then selectively fuse them. Figure 3(a) depicts an overview of our proposed method.
Figure 2: Comparison between previous datasets and our proposed MVG dataset. (a) is the dataset used by the previous work, which only have clothing and person in the frontal-view. In contrast, our dataset (b) offers images from five different views.
Figure 3: (a) Overview of MV-VTON. It encodes frontal and back view clothing into global features using the CLIP image encoder and extracts multi-scale local features through an additional encoder $\mathcal{E}_{l}$. Both features act as conditional inputs for the decoder of backbone. Besides, both features are selectively extracted through view-adaptive selection mechanism. (b) Soft-selection modulates the clothing features on frontal and back views, respectively, based on the similarity between the clothing’s pose and the person’s pose. Then the features from both views are concatenated in the channel dimension.
View-Adaptive Selection
@@
For multi-view virtual try-on task, given the substantial differences between the frontal and back views, as illustrated in Figure 2(b), it’s imperative to extract and assign the features of frontal and back view clothing for the person tendentiously. Actually, based on the pose of the target person, we can determine which view of clothing should be given more attention during the try-on process. For example, if the target pose resembles the pose in the fourth column of Figure 2(b), it’s evident that we should rely more on the characteristics of the back view clothing to generate the try-on result. Specifically, we propose a view-adaptive selection mechanism to achieve this purpose, including hard- and soft-selection.
@@
Hard-Selection for Global Clothing Features. We deploy a CLIP image encoder to extract global features of clothing. During this process, we perform hard-selection on the frontal and back view clothing based on the similarity between the garments’ pose and the person’s pose. It means that we only select one piece of clothing that is closest to the person’s pose as the input of the image encoder, since it is enough to cover global semantic information. When generating pre-warped clothing for $\mathcal{E}(a)$, the selection is also performed. Implementation details of hard-selection can be found in the supplementary material.
@@
Soft-Selection for Local Clothing Features. We utilize an additional encoder $\mathcal{E}_{l}$ to extract the multi-scale local features of frontal and back view clothing, which in the $i$-th scale are denoted as $c_{f}^{i}$ and $c_{b}^{i}$, respectively. When reconstructing the try-on results, it may be insufficient to rely solely on the clothing from either frontal or back view under certain specific scenes, such as the third column shown in Figure 2(b). In these cases, it may be necessary to incorporate clothing features from both views. However, simply combining the two may lead to confusion of features. Instead, we introduce soft-selection block to modulate their features, respectively, as shown in Figure 3(b).
First, the person’s pose $p_{h}$, frontal-view clothing’s pose $p_{f}$, and back view clothing’s pose $p_{b}$ are encoded by the pose encoder $\mathcal{E}_{p}$ to obtain their respective features $\mathcal{E}_{p}(p_{h})$, $\mathcal{E}_{p}(p_{f})$, and $\mathcal{E}_{p}(p_{b})$. Details of the pose encoder can be found in the supplementary material. When processing frontal-view clothing, in $i$-th soft-selection block, we map $\mathcal{E}_{p}(p_{h})$ and $\mathcal{E}_{p}(p_{f})$ to $P^{i}_{h}$ and $P^{i}_{f}$ through a linear layer with weights $W_{h}^{i}$ and $W_{f}^{i}$, respectively. We also map $c_{f}^{i}$ to $C^{i}_{f}$ through a linear layer with weights $W_{c}^{i}$. Then, we calculate the similarity between the person’s pose and frontal-view clothing’s pose to get the selection weights of frontal-view clothing, i.e.,
$weights=softmax(\frac{P_{h}^{i}(P_{f}^{i})^{T}}{\sqrt{d}}),$ (4)
@@
where $weights$ represents the selection weights of frontal-view clothing, and $d$ represents the dimension of these matrices. Assuming that the person’s pose is biased towards the front, as depicted in the second column of Figure 2(b), the similarity between the person’s pose and the front view clothing’s pose will be higher. Consequently, the corresponding clothing features will be enhanced by $weights$, and vice versa. The features of back view clothing $c_{b}^{i}$ undergo similar processing. Finally, the two selected clothing features are concatenated along the channel dimension as the local condition $c_{l}^{i}$ of backbone.
Figure 4: Overview of the proposed joint attention blocks.
Joint Attention Blocks
@@
Global clothing features $c_{g}$ provide identical conditions for blocks at each scale of U-Net, and multi-scale local clothing features $c_{l}$ allow for reconstructing more accurate details. We present joint attention blocks to align $c_{g}$ and $c_{l}$ with the current person features, as shown in Figure 4. To retain most of the semantic information in global features $c_{g}$, we use local features $c_{l}$ to refine some lost and erroneous detailed texture information in $c_{g}$ by selective fusion.